The annual profit for a firm depends upon the number of units produced. the function which describes the relationship between profit P (stated the number of units produced x is


P-0.01x²+5,000x - 25,000


(a) Determine the number of units x which will result in maximum profit. (b) What is the expected maximum profit?


1
Expert's answer
2021-05-24T16:18:00-0400

Given Profit-

"P=-0.01x\u00b2+5,000x - 25,000"


Differentiate P w.r.t. x-


"P'=-0.02x+5000"


Putting P'=0, we get-


"-0.02x+5000=0\\\Rightarrow x=\dfrac{5000}{0.02}=250000"


Again Differentiating P' w.r.t. x-


"P''=-0.05<0,"


P is maximum at x=250000


(a) The unit which results in the maximum profit is 250000.


(b) The expected maximum profit is


"P=-0.01(250000)^2+5000(250000)-25000\\=-625000000+1250000000-25000\\=624975000"